# Sorting-based FPGA Sliding Window Aggregation Engine without off-chip Memories

### v2

@@ 4.5. Resource utilisation
Table 1 introduces the look-up table (LUT) and flip-flop register (FF) utilisation as reported by Vivado 2024.1 (and 2020.1 for the unmodified sorter results for consistency with literature) for the selected target device. The presented utilisation is for the encapsulating AXI peripheral. Note that slight variations can occur when further exploring optimisation directives, interfacing, register placement etc., though they remain consistent between all presented runs.
Table 1. Vivado resource utilisation report summary
Design Group operators LUT FF
Sorter (Papaphilippou et al., 2020) N/A 34,101 47,122
Sorter & Group-by engine min., max., sum, count, distinct count 36,653 49,544
Adapted sorter1 & SWAG engine min./med./max. 41,963 54,437
Adapted sorter1 & SWAG engine min., max., sum, count, distinct count, min./med./max. 43,441 54,694
1with cardinality support, as introduced in section 3.3.
With respect to the group-by aggregate engine, we can see that there is about a 2.5K overhead on the number of Lookup-tables (LUTs) when becoming an add-on to an existing unmodified sorter. All engine designs here use a register per pipeline stage, which also introduced a 2.4K overhead on flip-flop (FF) registers for the engine.
The group-by engine essentially improves on the design of the PRRA-based distinct counter (Papaphilippou et al., 2023) of figure 1. If we compare its utilisation (2.55K LUT, 2.42K FF) to the unmodified PRRA (1.26K LUT, 2.29K FF (Papaphilippou et al., 2023)), it is easy to conclude that it uses at least 1.9x fewer FFs than any solution with two PRRAs, and similar LUTs before even adding the aggregation logic. This can be explained with the help of a simple model: when we abstract the node complexity to 1 for the butterfly network switches and all other entities, the complexity of the example PRRA-based engine of figure 1 decreases from $3P+2\times\mathrm{PRRA}$ to $12P+1\times\mathrm{PRRA}=2P\times\log_{2}(P)+P+$ 1 entities.
All changes to build the SWAG engine total around 7K LUTs and 5K FFs, and this includes the cardinality calculation support within the sorter for each tuple (see section 3.3). In terms of block RAM, it needs to accommodate $k^{2}$ tuples for the sliding window buffer and $k^{2}$ for the merge phase of the sorter, where $k$ is the sorter capacity. Since $k=128$, this totals $3k^{2}\times 64\textit{bits}=3\textit{Mibs}\approx 85\times 36\textit{Kibs}$, resulting in 89 BMEMs.
5. Related work
There is a moderate amount of research on database acceleration focusing on analytics (Papaphilippou and Luk, 2018). Similar to earlier works that accelerated joins (Casper and Olukotun, 2014; Owaida et al., 2017), there are tradeoffs when selecting between a sort and a hash-based solution (Moghaddamfar et al., 2023; Absalyamov et al., 2016), each with separate adaptation challenges (Mätas et al., 2022; Eryilmaz et al., 2021). For instance, even with a sorter-based pipeline, supporting arbitrary data in a join accelerator caused stalls and increased its memory requirements (Casper and Olukotun, 2014). This related to the existence of duplicates of unknown cardinality, which is somewhat reminiscent of median support, which affects the complexity of SWAG designs.
Group-by-aggregate operators have usually been a part of larger analytics designs (Lisa et al., 2018), as with the presented engine. This is due to the bandwidth limitations (Eryilmaz et al., 2021), which makes isolated solutions less practical. A group-by-aggregate-focused work demonstrated a speedup of 10x over the CPU (Absalyamov et al., 2016), but used multi-tasking, hashes and off-chip memory on a discontinued platform. A recent group-by-aggregator uses sorting to accelerate these operations (Xue et al., 2024), though it implements an SQL-based data management system, which is still inline with the observation on the scarcity of out-of-context group-by-aggregate accelerators.
On the other hand, SWAG engines exist stand-alone, also making direct comparisons more insightful. A state-of-the-art FPGA solution provides a notable speedup over prior work, but still operates with no ability to process multiple elements per cycle (Geethakumari and Sourdis, 2021b, 2023). It involves the implementation of a multi-level memory hierarchy based on off-chip memories (Geethakumari and Sourdis, 2021a), which is reflected on the complexity of the design. In contrast, our proposed SWAG engine is able to be evaluated on a resource-constrained embedded platform, also without the need for compression (Geethakumari and Sourdis, 2021b), while still having significant advantages.
Table 2. SWAG engine implementation comparison, inspired by (Geethakumari and Sourdis, 2023)
DFE(Ramakrishnan Geethakumari et al., 2019) MLQ(Geethakumari and Sourdis, 2021a) StreamZip(Geethakumari and Sourdis, 2021b, 2023) Proposed work
Group operators min./med./avg/max. min. min., max., sum, count, dist. count, min./med./max.
@@ 5. Related work
[…]
1Intel’s ALM is not directly comparable, but implies more logic per unit.
2All operators except summation also support IEEE-754 floating points.
[…]
6. Discussion and future work
The isolated task of group-by-aggregate (non-SWAG) is not necessarily expensive on other platforms, such as when using hash tables for smaller arrays instead of sorting on CPUs. Still, the engine simplifies the enablement of this functionality in programmable logic, such as for SWAG and high-throughput pipelined database analytics accelerators (Sun et al., 2021) to reduceunnecessary data transfers (Lisa et al., 2018).
The engines are highly customisable and parameterisable. The selection mechanism inside the SWAG engine can easily be generalised to support a percentile operator to select an arbitrary number of percentiles based on the pre-calculated cardinality. The average operation is supported in simulation, but is omitted for brevity, in order to avoid shifting the focus to floating point division optimisations. The throughput and latency limitations mentioned for SWAG in section 4.1 can be addressed. When a higher throughput is needed for large windows, a wider $P$ can be used, or two sorters can be combined and de-multiplexed using rate converters. A lower latency can be achieved by using lower $k$ values for $\textit{WS}\leq k$, or a higher $P$ for $\textit{WS}>k$. If the missing $\textit{WS}=256$ mode is needed, a build with $k=64$ or $256$ would suffice. Alternatively, this inherited limitation of the utilised sorter can be addressed at its source, to enable splitting the sorted sublists into smaller pieces to also conserve the merging throughput.
Other future work includes the automation of synthesising customised aggregation engines, such as by developing a high-level synthesis (HLS) library. Additional effort could be put into automating the inference of such circuits and modular components. The proposed aggregation engine could also be explored as a functional unit inside CPUs to work as a set of dedicated SIMD instructions.
7. Conclusions
The increasing memory datapath widths are becoming increasingly challenging to exploit with existing programming techniques and models. This includes naturally parallel problems such as aggregation. A modular and highly-adaptable engine is presented that uses prefix-scan and reverse butterfly topologies to perform a wide range of aggregation tasks with high-throughput as a non-blocking pipeline. It is then adapted to work as a highly competitive sliding window aggregation (SWAG) engine by incorporating a sorter that sorts each incoming buffered window. A wide selection of additions are proposed to an existing high-throughput sorter to append group cardinality information to the incoming tuples, in order to facilitate the calculation of the median operator. The resulting designs are shown to have a low hardware overhead, while supporting a superset of the functionality of alternatives. The SWAG engine exhibits up to hundreds of times of speedup when compared to a CPU-only approach on our embedded evaluation platform, and is able to process multiples of the input rate of the state-of-the-art on a datacenter grade FPGA. This is achieved without the need for off-chip memories.

### v3

@@ IV-F Resource utilisation
Table I introduces a summary of the evaluated engines alongside their look-up table (LUT) and flip-flop register (FF) utilisation as reported by Vivado. The presented utilisation is for the corresponding encapsulating AXI peripheral. Note that slight variations can occur when further exploring optimisation directives, interfacing, register placement etc., though they remain consistent between all presented runs.
TABLE I: Vivado resource utilisation report summary
Engine, input widths Sorter, input widths Group operators LUT $F$
@@
1Also appending cardinality information.
With respect to the group-by aggregate engine, we can see that there is about a 2.5K overhead on the number of Lookup-tables (LUTs) when becoming an add-on to an existing unmodified sorter. All engine designs here use a register per pipeline stage, which also group-by introduced a 2.4K overhead on flip-flop (FF) registers for this engine.
The SWAG engine without group IDs consumes the least amount of resources in table I. This is because of the absence of groups, which yields less resources for sorting, as it uses 32-bit values only. In terms of percentage of the ZU3 device, the LUT and FF utilisation is 42% 30% respectively. For all sorters/cardinality counters in the designs, the tested internal sorter capacity is steady at $k=128$. This is despite being the only one with $P=8$ to saturate the datapath, processing 8 $\times$ 32-bit values per FPGA cycle.
For SWAG with groups and the $min./med./max.$ operator, if we abstract all the SWAG functionality as an add-on to the sorter, the overhead totals around 8K LUTs and 7K FFs. These values includes the cardinality calculation support within the sorter for each tuple of section III-D, which makes this new hybrid sorter-cardinality counter an integral part of the design.
@@
IV-G Scalability
The proposed engines are scalable to high degrees of parallelism $P$ and are able to saturate the bandwidth for current and future systems with a wide datapath. The total amount of data that can be produced or consumed by the engine per cycle is $P$ times the data width. An exploration of the impact of $P$ and the data width on the resources is conducted through out-of-context synthesis. The tuple (data) width is halved and used by the group ID and key fields for simplicity. The synthesis is done with Yosys [21] for a more generic but modern LUT6 architecture, which can serve as high-level guidelines for the resource utilisation expectations for a variety of platforms.
Figure 13: Synthesis results when varying tuple width and $P$.
Figure 13 summarises the results of this exploration for the SWAG engine without the accompanying sliding window buffering and sorter. The sliding window buffering logic is simplistic, hence it is omitted from the scalability study, as it mainly consists of block RAM. Since our modifications to the sorter are relatively lightweight, the readers can also refer to the sorter’s original description on its scalability [15]. For the ranges of interest, the registers and LUTs scale almost linearly with $P$ and the tuple width. The fastest growing component is the reverse butterfly network, with a complexity of $\Theta(P\log_{2}(P))$ switches (similar for the prefix scans), though the slight upward trend is not observed in these ranges yet. Our use case has yielded a low enough register utilisation for our purposes, but future work can also focus on retiming to considerably reduce the register utilisation when necessary.
Overall, $P$ seems more influential to the LUT and register use than the tuple width, as it scales more rapidly for the studied ranges. Even the highest points here ($P=64$, tuple width 256 bits) are feasible with today’s datacenter FPGAs, though any complications would first relate to how easy it is to establish a sorted input stream with such a high throughput.
This scalability study focuses on the aggregation part for SWAG with groups without loss of generality, since it is the most feature-rich version of the base pipeline. The obtained curves for the other pipelines are very similar, but they are omitted for brevity.
IV-H Throughput and latency
Aggregation engines are often found as streaming accelerators, hence it can be a requirement to be able to cope with the line rate of the system, especially when the stream is continuous. As a non-blocking pipeline, the main aggregation pipeline supports the full throughput of $P$ elements per cycle without any backpressure for tasks such as group-by. Its pipeline latency of $1+2\log_{2}P$ cycles is negligible for streaming data, which could be minimised further with register retiming.
When it is extended to implement the SWAG engine, it becomes a blocking design (i.e. its “ready” signal or equivalent is not always high). The amount of backpressure it applies on its producer mostly relates to the window advance (WA) in relation to its window size (WS). This is because, when $\mathrm{\textit{WA}}<\mathrm{\textit{WS}}$ it has the potential to produce more output than input [22]. Specifically, assuming that other than the sliding window buffering there is no other backpressure from the rest of the engine or system, the highest throughput that can be achieved by an ideal implementation is $P\mathrm{\times\textit{WS}}/\mathrm{\textit{WA}}$, where $P$ is the degree of parallelism.
Figure 14: Final input rate influenced by backpressure, for $P=4$.
Figure 14 provides simulation results using Icarus Verilog [23] that show the input rate by counting the number of input tuples per cycle for a relatively long input size and $P=4$. On the left plot, it can be observed that the input rate peaks near 4, especially when $\mathrm{\textit{WS}}=\mathrm{\textit{WA}}=k=128$. This is since the sorter processes the data as they arrive, there is no buffer “replay” for $\mathrm{\textit{WS}}=\mathrm{\textit{WA}}$, and any other buffering within the design such as the PMT is comparatively diminishing. After $\mathrm{\textit{WS}}=256$, the maximum input rate is up to half of the available, since the sorter needs an internal second pass for the merging phase. On the right plot, the numbers are divided by $\mathrm{P\times\textit{WS}}/\mathrm{\textit{WA}}$ to uncover any inefficiencies, with the main observation being the halving of the throughput when $\mathrm{\textit{WS}}>k$.
The plots of figure 14 have a triangular shape, as the supported window advance cannot be greater than the window size, i.e. $\mathrm{\textit{WA}}\leq\mathrm{\textit{WS}}$. At $\mathrm{\textit{WS}}=256$, there are no results due to a limitation within the sorter (merging at least $P$ lists of $k$ elements). If the missing $\textit{WS}=256$ mode is needed, a build with $k=64$ or $256$ would suffice. Alternatively, this inherited limitation can be addressed at its source [15], to enable splitting the sorted sublists into smaller pieces.
Figure 15: Simulated latency for $P=4$, $k=128$ and $f_{\textit{target}}=125$ MHz
Figure 15 presents a similar experiment that measures the latency between the time the first tuple enters into the engine and the time of the first output. As shown, the WS does not influence the latency up to $\textit{WS}=128$, since the linear sorters in the sorter have a fixed latency ($k/P$) for an input to be extracted. For $\textit{WS}>k$ the latency scales linearly with WS, as the sorted windows are merged with a fixed rate $P$, and the merge phase is compulsory. WA does not impact the latency.
These results are generalisable to all queries using the SWAG architecture and tuple widths, because the output format does not influence the backpressure as long as up to $P$ tuples are accepted at the output of the engine.
V Related work
There is a moderate amount of research on database acceleration focusing on analytics [24]. Similar to earlier works that accelerated joins [25, 2], there are tradeoffs when selecting between a sort and an indexing-based solution [26, 27], each with different adaptation challenges [28, 29]. For instance, even with a sorter-based pipeline, supporting arbitrary data in a join accelerator caused stalls and increased its memory requirements [25]. This relates to the existence of duplicates of unknown cardinality, which is somewhat reminiscent of median support, which affects the complexity of SWAG designs.
TABLE II: SWAG engine implementation comparison
DFE [30] StreamZip [7] Presented engine implementations Mueller et al. [31] Oge et al. [32]
SWAG context w/ groups (window per group) w/ groups (single window) w/o groups time-based
[2pt/1pt] Group operators min., med., avg, max. min. min., max., sum, count, dist. count, min./med./max. med. min., max., sum, count
[2pt/1pt] Max. input rate 0.5 4 8 1
[2pt/1pt] Line rate 128-bit@78.13 MHz 128-bit@250 MHz 32-bit@100 MHz 128-bit@200 MHz
[2pt/1pt] FPGA device Intel 5SGXAB AMD ZU3EG XC2VP30 XC6VLX240T
[2pt/1pt] Logic (K) 861 1041 2191 3081 43 LUT, 55 FF 29 LUT, 42 FF 6.4 LUT, 3.8 FF <<1.5 LUT, <<3 $F$
[2pt/1pt] BRAM (M20K) 1136 1584 1796 2139 89 RAMB36E2 92.5 RAMB36E2 0 17 RAMB36
[2pt/1pt] BRAM (Mbits) 22.4 31.2 35.4 42.1 3.3 3.5 0 0.6
[2pt/1pt] DSPs 0 165 0
@@ V Related work
[2pt/1pt] Key format fixed-point fl.-point fl.-point (lossy) integer, fl.-point2 integer
@@
[2pt/1pt] WA range {1,2,4,,4K1,2,4,...,4K} $\{4,8,...,16K$} $\{8,16,...,16K$} {1} N/A
[2pt/1pt] Off-chip memory DRAM DRAM & 72 MB QDR-SRAM 0
1Intel’s ALM is not directly comparable, but implies more logic per unit. 2All operators except summation also support IEEE-754 floating points.
3For WS and WA here, $K$ denotes kibi (1024), not kilo (1000).
V-A Group-by aggregation
Group-by aggregate operators have usually been a part of larger analytics designs [33]. This is due to the bandwidth limitations [29], as with co-processors’ reduced performance when it comes to accessing main memory. This makes stand-alone solutions more challenging to become practical.
A group-by-aggregate-focused work demonstrated a speedup of 10x over the CPU [27], but used multi-tasking, hashes and off-chip memory on a discontinued platform. A recent group-by-aggregator uses sorting to accelerate these operations [34], though it implements an SQL-based data management system as a whole, which is still inline with the observation on the scarcity of out-of-context group-by-aggregate accelerators.
Despite this, the acceleration of specific group-by queries is useful in certain industrial contexts, such as distinct count per group for customer analytics [15]. This operator is supported by the presented engine, but can also be implemented by combining two PRRAs for intermediate sparsity removal, as shown in figure 16.
Figure 16: Modular design of the distinct count query (right) [15] with example data in orange, based on the PRRA implementation (left) [4].
When we compare the engine’s utilisation (2.55K LUT, 2.42K FF, see table I) to the unmodified PRRA (1.26K LUT, 2.29K FF [4]), it is easy to conclude that it uses at least 1.9x fewer FFs than any solution with two PRRAs, and similar LUTs. This can be explained with the help of a simple model: when we abstract the node complexity to 1 for the butterfly network switches and all other entities, the complexity of the example PRRA-based engine of figure 16 decreases from $3P+2\times\mathrm{PRRA}$ to $2P+1\times\mathrm{PRRA}=2P\times\log_{2}(P)+P+1$ entities. This is while supporting multiple operators reconfigurable at runtime. In this instance, our proposed group-by engine reduces the best known resource complexity for performing distinct count per group, while being highly adaptable for other such advanced operators.
V-B SWAG engines
On the other hand, SWAG engines exist stand-alone, making comparisons more direct. Table II presents a comprehensive quantitative comparison on how the proposed approach compares to the state-of-the-art and representative works.
Starting from the left of table II, the comparison begins with SWAG with groups. State-of-the-art FPGA solutions provide a notable speedup over prior work [30], but still operate on up to a single tuple per cycle [7, 35]. They often involve the implementation of a multi-level memory hierarchy based on off-chip memories, which is reflected in the complexity of the design. In contrast, our proposed engine achieves to be evaluated on a resource-constrained embedded platform (no loss of generality to high-end FPGAs).
An important difference between the presented SWAG engine and the selection of competing engines for SWAG with groups is that there is a single sliding window, as opposed to a window per group (elaborated in section VI). In this case, our design uses no DRAM and a fraction of the BRAM at 3.3 versus 22.4 Mbits for the most lightweight of the alternatives [30]. The competing engines have a target frequency of 150 MHz, but at 125 MHz internally, it is able to saturate a line rate over 3 times higher. In terms of throughput, the competing variations in the table are essentially exchanging resources and functionality for performance (not shown). For example, DFE utilises around $1/7$ of the input rate, even when WS==WA. The compression in StreamZip [7] causes the throughput to scale differently according to the relationship between WS and WA, which can be beneficial in certain applications with a low WA.
@@ V-B SWAG engines
When it comes to SWAG without groups, Mueller et al. propose a sliding median operator, but was more of a proof of a generalisable concept, also given the state of the technology of the day [31]. The sliding window size is fixed to only 8, hence the reduced need for BRAM and DRAM. The median calculation is based on shift registers and a sorting network, which, even with newer technology, do not have the scalability of more sophisticated designs. For instance, the merge-capable linear sorters [15] enable a practical use with WS of thousands of elements in the presented engines. It operates at a single input per cycle, though parallelism is demonstrated for multiple workers operating on independent streams. A single worker has a throughput of 100 MT/s, while our approach reaches 1 GT/s.
Finally, a time-based SWAG engine is indicatively compared with the proposed engines due to its similarity in certain aspects [32]. While groups are supported in its input similar to count-based SWAG with groups, the results are only for a single group. It is the most space efficient design, but lacks the high throughput and more advanced operators. It focuses on achieving a steady latency and throughput for different WS and WA ratios, but each of those combinations requires a separate bitstream, since it is not runtime configurable. At a single tuple per cycle, its throughput is 200 MT/s.
Overall, our SWAG engine’s contributions stand between SWAG with groups and without groups including, enabling big data processing with more advanced operators for simple devices. It is worth noting that the comparison is, hence, relatively diverse, as it features partly different systems and operator combinations, and further customisation is possible.
VI Discussion and Future work
The engines are highly adaptable and parameterisable.
Other operators
The selection of operators in the aggregation pipeline can easily incorporate other associative operators such as multiplication, bitwise exclusive-OR etc. This can be achieved by including the corresponding logic inside the $n$ and $n^{\prime}$ entities. The average operation is not explored beyond simulation to avoid shifting the focus to floating point division optimisation. This is without loss of generality, only requiring $P$ dividers, usually consuming a few hundred LUT/FFs according to the use case, and without a perceptible timing overhead [36].
Similarly, other types of averaging are possible by combining simultaneous associative operators together at the $n^{\prime}$ stage. As an example, combining multiplication and count with a root circuit, can provide the geometric mean. Another instance is with the selection mechanism inside the SWAG engine, which can be generalised to support the percentile operator to select arbitrary percentiles based on the pre-calculated cardinality. Other selection queries could use more specialised criteria. Future applications could exploit the additional information in the engine, such as the order of the sorted data, which has already been used in the distinct count algorithm (step $(c)$ in section II-B) to essentially make it an associative operation.
SWAG context
In the comparison of section V-B, a unique aspect for the SWAG engine with groups is the single window, while still providing per group statistics. Per-group window SWAG, although supported, might be less trivial to efficiently exploit the presented architecture, as the sorting advantage might diminish when “recycling” more data on each window advance. Thus, it could be seen as an approximation for when the task involves a higher number of groups and there is a need for multiple windows. In practice, keeping live statistics for all groups may not be necessary [32], and a simple filtering step to isolate the group IDs would still dedicate the whole window to the desired group.
Yet, the notion of having a common window for grouped data is widely useful for other SWAG tasks. One such case is for supporting more complex SWAG queries without groups, where the group field can carry the value instead, for operating in the context of the values. This includes the mincount or maxcount operator (i.e. counting the instances of the minimum or maximum value within the window) [37], which can be implemented as a group-by query where the count operator is only activated on the first or last group, corresponding the minimum or maximum respectively.
It is also especially useful when using a time-based window eviction policy, such as when summarising statistics for the last few seconds [32]. The presented aggregation pipeline, as well as the sorter extensions for cardinality calculation actually support both count-based and time-based SWAG, since each cell automatically carries an oscillating counter to distinguish between arbitrarily-sized windows, assuming WS is sufficient. This essentially means that the sliding-window mechanism can be adjusted accordingly for the whole engine to support different window contexts. Time-based aggregation requires minimal modification at step $(a)$ in section III-B to produce windows based on a timestamp. The present evaluation can be considered representative for time-based SWAG, if the order of the data has a time notion, such as when each new tuple is generated on a fixed interval.
Specialisation and optimisation
The presented engines and base pipeline can be simplified further according to the desired operator selection. If only associative operators (e.g. sum) are needed, then the requirement for the sorted input is only with respect to the group. For SWAG, this essentially means that the comparison width of the small-scale sorter can be constrained to that of the group to minimise its complexity. For non-SWAG, this would also be reflected in the accompanying sorter, or in software equivalents that provide the sorted stream. If selection operators such as the median are not present, then the cardinality logic (section III-D) is not needed. Hence, the disjoint prefix-scan steps can also be merged back. In this scenario, alternative sorters would also suffice for SWAG. When this holds and the input does not carry group information, then a sorter is not needed at all.
As demonstrated in the study of section IV-F are scalable for wide datapaths. The throughput and latency behaviour of the SWAG engine can be optimised further. When a higher throughput is needed for large windows, a wider $P$ can be used, or two sorters can be combined and de-multiplexed using rate converters. A lower latency can be achieved by shortening the linear sorter modules, such as by using lower $k$ values for $\textit{WS}\leq k$, or a higher $P$ for $\textit{WS}>k$.
Other future work includes the automation of synthesising customised aggregation engines, such with a high-level synthesis (HLS) library. It could also be explored as a functional unit inside CPUs to work as dedicated SIMD instructions [38].
VII Conclusions
The increasing memory datapath widths are becoming increasingly challenging to exploit with existing programming techniques and models. This includes aggregation, a naturally parallel problem. A modular and highly-adaptable aggregation pipeline is presented that uses prefix-scan and reverse butterfly topologies to perform a wide range of aggregation tasks with high-throughput. The current design improves the theoretical and synthesis-based circuit complexity for specialised group-by applications such as for counting unique elements per group. It is then adapted to work as a competitive sliding window aggregation (SWAG) engine by incorporating a modified sorter that sorts each incoming buffered window on chip. The sorter includes lightweight algorithmic modifications that append group cardinality information to the incoming tuples to assist with median calculation. The resulting designs are shown to have low hardware overheads, while often supporting a superset of the functionality of alternatives. The SWAG engine exhibits up to hundreds of times of speedup when compared to CPU-only approaches on our embedded evaluation platform, and is able to process multiples of the input rate of the state-of-the-art on datacenter-grade FPGAs. This is achieved with the practical assumption of a single window for all group IDs, which in turn has a wider applicability for SWAG workloads. The highly-optimised on-chip pipeline eliminates the need for off-chip memories such as DRAM, and reduces the latency for relatively complex aggregation tasks.
Acknowledgements
The support of UK EPSRC (grant number UKRI256, EP/V028251/1, EP/N031768/1, EP/S030069/1, and EP/X036006/1) is gratefully acknowledged. Prof David Gregg has provided valuable feedback on an early draft.